Find two consecutive integers whose product is 72?
Answer by nerdybill(7384)

Find two consecutive integers whose product is 72?
.
Let x = first integer
then
x+1 = consecutive integer
.
x(x+1) = 72
x^2+x = 72
x^2+x-72 = 0
(x+9)(x-8) = 0
x = {-9,8}
.
Answer: 8 and 9